# Improving Multimodal Learning Balance and Sufficiency through Data Remixing

### v1

@@ 1 Introduction
Multimodal learning (ngiam2011multimodal) is a rapidly evolving field in artificial intelligence, aimed at enhancing the perception and decision-making capabilities of models by integrating data from diverse modalities, including vision, sound, and text (vision+). However, existing multimodal learning methods often face challenges in fully integrating rich multimodal knowledge across different modalities. Due to the inherent differences in data representation and distribution across modalities, their optimization trajectories differ significantly, resulting in imbalanced learning when multiple modalities are jointly trained under a unified objective.
@@
Specifically, multimodal models often prioritize learning the most discriminative features from the strong modality, suppressing the training of the weak modality and causing it to become lazy, known as modality laziness (wang2020makes; huang2022modality; du2021improving).
Several methods have been proposed to address these issues. Some focus on alleviating modality laziness by enforcing the weak modality through modality-specific optimization objectives, such as adjusting task-specific supervision (wang2020makes; xu2023mmcosine; du2023uni), introducing prototype learning (fan2023pmr), or using knowledge distillation (du2021improving). Other methods aim to align optimization speeds by modifying learning rates (sun2021learning) and gradients (peng2022balanced; li2023boosting; sun2023graph; fu2023multimodal; kontras2024improving; wei2024mmpareto) based on unimodal performance. In contrast, some approaches attempt to regulate modality inputs to more explicitly enhance the learning of weak modalities, such as by augmenting weak modality samples (greedy; wei2024enhancing) and masking strong ones (zhou2023adaptive), or by decoupling multimodal learning into unimodal tasks to avoid modality laziness (mla). However, most of the existing methods align the optimization speeds of different modalities to alleviate modality laziness, without addressing the fact that even when modality balance is achieved, differing optimization paths can still cause interference during modality optimization. In multimodal learning, the gradient update directions of different modalities are inconsistent (fan2023pmr), leading to cross-modal interference during batch gradient descent (Figure 3), which can be called modality clash. This inconsistency causes the modalities to deviate from their expected optimization paths, as shown in Figure 1, resulting in insufficient learning across all modalities and limiting the effectiveness of multimodal learning. Moreover, some methods are constrained by specific model architectures (he2022multimodal; lin2023variational; mla) or hinder training efficiency (wei2024enhancing; mla), which further limits their applicability.
@@ 2.1 Balance with Multimodal Joint Input
wang2020makes; du2021improving discover that the optimization speed varies across different modalities. Therefore, when optimizing a multimodal model with a unified objective, the weak modality fails to learn adequately, leading to modality laziness. Most methods maintain multimodal data inputs during training, achieving modality balance by adjusting optimization objectives and speeds.
Some try to alleviate modality laziness by enforcing the weak modality through modality-specific optimization objectives wang2020makes; du2021improving; xu2023mmcosine; du2023uni; fan2023pmr. wang2020makes propose Gradient-Blending, which calculates the optimal mixing mode of modality losses by determining the overfitting situation for each modality. du2021improving attempt to distill knowledge from well-trained unimodal models to enhance the unimodal encoders. fan2023pmr introduce the prototype cross-entropy loss for each modality to accelerate the slow-learning modality. Others try to align optimization speed by changing learning rates (sun2021learning) or modifying gradients (peng2022balanced; li2023boosting; sun2023graph; fu2023multimodal; kontras2024improving) according to unimodal performance. sun2021learning dynamically adjust the learning rates of different modalities based on the unimodal predictive loss. peng2022balanced adaptively modulate the gradients of each modality by monitoring the discrepancy in their contribution to the learning objective. These methods promote modality balance and enhance the expressive power of multimodal models. However, they fail to recognize that even when modality laziness is addressed, modality clash persists, meaning that multimodal capability remains limited.
@@ 2.2 Balance with Unimodal Single Input
Some methods attempt to use alternating or selective unimodal inputs to promote multimodal learning, thereby avoiding interference from other modalities and enhancing the expressive power of multimodal models (greedy; mla; wei2024enhancing; zhou2023adaptive). greedy measure the relative speed of each modality and train unimodal branches to fully utilize them. mla decompose multimodal learning into alternating unimodal learning with gradient modification to preserve cross-modal interactions. wei2024enhancing design a sample-level modality contribution evaluation based on Shapley values and perform resampling of the weak modality to enhance specific modality training. These methods, which rely on unimodal inputs, seem to avoid modality clash. However, upon closer inspection, we find that modality clash is introduced at the batch level. Like the methods in Section 2.1, they have not analyzed or addressed the root cause of modality clash.
Overall, current work focuses on addressing modality laziness but fails to recognize that modality clash still exists in balanced multimodal learning. Meanwhile, some methods are constrained by specific model architectures (greedy; mla) or hinder training efficiency (mla; wei2024enhancing), limiting their applicability. In this paper, we aim to simultaneously mitigate modality laziness and modality clash while ensuring compatibility with various fusion methods and model architectures.
@@ 3.1 Model formulation
In this paper, we focus on the multimodal discrimination task, following related works (peng2022balanced; fan2023pmr; wei2024enhancing; mla). For convenience, we consider two input modalities: $m_{a}$ and $m_{v}$. The training dataset is $\mathcal{D}=\{x_{i},y_{i}\}_{i=1,2,\ldots,N}$. Each $x_{i}$ consists of multimodal inputs, i.e., $x_{i}=(x_{i}^{a},x_{i}^{v})$. $y_{i}\in\{1,2,\ldots,M\}$, where $M$ is the number of classes. The multimodal model is trained using batch gradient descent, with data $\boldsymbol{x}=(\boldsymbol{x}^{a},\boldsymbol{x}^{v})$ sampled from a batch $B$.
We use a multimodal model consisting of two unimodal branches for prediction. Each branch has a unimodal encoder, denoted as $\phi^{a}$ and $\phi^{v}$, used to extract features from the corresponding modality of $\boldsymbol{x}$. The encoder outputs are represented as $\boldsymbol{z}^{a}=\phi^{a}(\theta^{a},\boldsymbol{x}^{a})$ and $\boldsymbol{z}^{v}=\phi^{v}(\theta^{v},\boldsymbol{x}^{v})$, where $\theta^{a}$ and $\theta^{v}$ are the parameters of the encoders. The results of the two unimodal encoders are fused in some way (concat; dicisionfuse) to obtain the multimodal output. We use cross-entropy (CE) loss as the loss function. To achieve better representation ability in the warm-up stage, we add a separate classification head for each modality, as wang2020makes do, and modify the loss function as follows:
@@ 3.2 Data Remixing Method
Decouple Multimodal Data. In multimodal learning, the model tends to fit the modality that optimizes faster, often converging before the other modality has been sufficiently learned (du2021improving). Therefore, appropriately reducing the training speed of the strong modality may be an effective strategy for balancing learning (peng2022balanced). However, modality-level control over the optimization process or objectives overlooks the imbalance at the sample level between modalities (wei2024enhancing), which limits the effectiveness of the balancing. Based on this premise, we consider promoting balance by assigning specific samples (in terms of difficulty and quantity) to different modalities, achieving a more precise and convenient modality balance.
We first define a sample-level unimodal capability evaluation method to achieve either-or data allocation, instead of using Shapley values as proposed by wei2024enhancing [^1] For $x_{i}$, if the unimodal predictions are correct with probabilities of 0.9 and 0.6, the Shapley values for both modalities would be the same, but the representational abilities are different., to avoid dataset expansion. For each sample $x_{i}$, we obtain unimodal prediction probabilities $p_{i}^{k}$, $k\in\{a,v\}$. The representational ability of each modality is then assessed by calculating the KL divergence between $p_{i}^{k}$ and a uniform distribution as
@@
where $p^{k}_{i,j}$ represents the predicted probability for class $j$, and $u_{j}=\frac{1}{M}$ corresponds to the uniform distribution. KL divergence measures the difference between two probability distributions. In the current application scenario, a smaller KL divergence indicates that the output is closer to a uniform distribution, meaning that the separability of the corresponding unimodal output is worse. This can be interpreted as insufficient training for the current modality on this sample. Therefore, we decouple the multimodal inputs for each sample and only retain the modality that performs the worst, masking input from other modalities to achieve accurate unimodal training (zhang2024multimodal; wei2024enhancing). That is, the new training set is as follows:
@@
Without loss of generality, we assume that the multimodal fusion method selected for the model is Concatenation, as done by others (fan2023pmr; peng2022balanced; wei2024enhancing). Let $\boldsymbol{W}\in\mathbb{R}^{M\times(d_{z^{a}}+d_{z^{v}})}$ and $\boldsymbol{b}\in\mathbb{R}^{M}$ represent the parameters of the linear classifier that produces the logits output. We can then express the output of the multimodal model (without softmax) as
@@ 4.1 Dataset and Experimental settings
CREMA-D (cremad) is an audiovisual dataset for emotion recognition, consisting of 7,442 video clips from 91 actors. The dataset includes six of the most common emotions: anger, disgust, fear, happy, neutral and sad. A total of 2,443 participants rated each clip for emotion and emotional intensity using three modalities: audiovisual, video only, and audio only. The entire dataset is randomly divided into a training and validation set of 6,698 samples and a test set of 744 samples, with a ratio of approximately 9:1.
Kinetic-Sounds (ks) is a dataset derived from the Kinetics dataset (kinetic), which includes 400 action classes based on YouTube videos. Kinetic-Sounds consists of 31 action categories, selected for their potential to be represented both visually and aurally, such as playing various instruments. Each video is manually annotated for human actions using Mechanical Turk and is cropped to 10 seconds, focusing on the action itself. The dataset comprises 19k 10-second video clips, split into 15k for training, 1.9k for validation, and 1.9k for testing.
Experimental settings. Unless otherwise specified, all feature extraction networks used in the experiments are ResNet-18 (resnet), trained from scratch. During training, we use the Adam (kingma2014adam) optimizer with $\beta=(0.9,0.999)$ and set the learning rate to $5\times 10^{-5}$. We obtain unimodal prediction results from the unimodal classification heads and also present the results of our method by masking specific modalities to zero, as in dropout and wei2024enhancing. All reported results are averages from three random seeds, with all models trained on two NVIDIA RTX 3090 GPUs using a batch size of 64.
@@ 4.2 Comparison with conventional fusion methods
We first compare our method with several representative multimodal fusion methods commonly used in deep learning frameworks: Concatenation (Concat) (concat), Summation (Sum), Decision Fusion (Decision) (dicisionfuse), FiLM (film), and Bi-Gated (gated). The results are shown in Table 1. We apply Data Remixing in combination with Concatenation as the representative fusion method for our approach. It is evident that our method significantly improves model performance across different datasets, with an increase of 8.20% on CREMAD and 5.40% on Kinetic-Sounds, outperforming other conventional fusion strategies. To further demonstrate the generalizability of our method, we combine Data Remixing with Summation and Decision Fusion, achieving substantial improvements across both datasets.
@@ 4.3 Comparison with imbalanced multimodal learning methods
Our method is primarily designed to address the issues of modality imbalance and insufficiency in multimodal learning. To evaluate the improvements, we compare our approach with several representative methods that target multimodal balance and sufficiency: G-Blend (wang2020makes), OGM-GE (peng2022balanced), Greedy (greedy), PMR (fan2023pmr), MLA (mla) and Resample (wei2024enhancing). For a fair comparison, we adopt Concatenation as the fusion strategy for the baseline, following the approach used in the above methods. The results, including accuracy (Acc) and dataset expansion factor (Factor), are presented for all methods.
@@ 4.4 Combination with complex cross-modal architectures
The above methods and experiments are conducted using simple fusion methods, where multimodal fusion occurs after the unimodal encoders or classifiers. To validate the applicability of the Data Remixing method in more complex multimodal scenarios, we combine it with two intermediate fusion methods: MMTM (joze2020mmtm) and CentralNet (vielzeuf2018centralnet). MMTM recalibrates the channel features of different CNN streams through squeezing and multimodal excitation steps, while CentralNet uses unimodal hidden representations alongside a central joint representation at each layer, performing fusion through a weighted summation learned during training.
@@ 4.5.2 Study of Decouple Multimodal Data
In the process of conducting sample-level evaluation and decoupling multimodal data, we choose to mask the better-performing modality and train with specific samples to alleviate modality imbalance. Figure 3 shows the distribution of samples assigned to different modalities during training. It is clear that the strong modality consistently receives fewer samples compared to the weak modality. Figure 3 illustrates the change in imbalance ratio $\rho$ (peng2022balanced) before and after decoupling the modality inputs. The results indicate that our method alleviates the modality imbalance, supporting our hypothesis that balancing modality representation via sample quantity and difficulty can improve model performance. Note that our method is applied after a 10-epoch warm-up stage.

### v2

@@ 1 Introduction
Multimodal learning (Ngiam et al., 2011) is a rapidly evolving field in artificial intelligence, aimed at enhancing the perception and decision-making capabilities of models by integrating data from diverse modalities, including vision, sound, and text (Zhu et al., 2024). However, existing multimodal learning methods often face challenges in fully integrating rich multimodal knowledge across different modalities. Due to the inherent differences in data representation and distribution across modalities, their optimization trajectories differ significantly, resulting in imbalanced learning when multiple modalities are jointly trained under a unified objective.
@@
Specifically, multimodal models often prioritize learning the most discriminative features from the strong modality, suppressing the training of the weak modality and causing it to become lazy, known as modality laziness (Wang et al., 2020; Huang et al., 2022; Du et al., 2021).
Several methods have been proposed to address these issues. Some focus on alleviating modality laziness by enforcing the weak modality through modality-specific optimization objectives, such as adjusting task-specific supervision (Wang et al., 2020; Xu et al., 2023; Du et al., 2023), introducing prototype learning (Fan et al., 2023), or using knowledge distillation (Du et al., 2021). Other methods aim to align optimization speeds by modifying learning rates (Sun et al., 2021) and gradients (Peng et al., 2022; Li et al., 2023; Sun et al., 2023; Fu et al., 2023; Kontras et al., 2024; Wei & Hu, 2024) based on unimodal performance. In contrast, some approaches attempt to regulate modality inputs to more explicitly enhance the learning of weak modalities, such as by augmenting weak modality samples (Wu et al., 2022; Wei et al., 2024) and masking strong ones (Zhou et al., 2023), or by decoupling multimodal learning into unimodal tasks to avoid modality laziness (Zhang et al., 2024b). However, most of the existing methods align the optimization speeds of different modalities to alleviate modality laziness, without addressing the fact that even when modality balance is achieved, differing optimization paths can still cause interference during modality optimization. In multimodal learning, the gradient update directions of different modalities are inconsistent (Fan et al., 2023), leading to cross-modal interference during batch gradient descent (Figure 3), which can be called modality clash. This inconsistency causes the modalities to deviate from their expected optimization paths, as shown in Figure 1, resulting in insufficient learning across all modalities and limiting the effectiveness of multimodal learning. Moreover, some methods are constrained by specific model architectures (He et al., 2022; Lin et al., 2023; Zhang et al., 2024b) or hinder training efficiency (Wei et al., 2024; Zhang et al., 2024b), which further limits their applicability.
@@ 2.1 Balance with Multimodal Joint Input
Wang et al. (2020); Du et al. (2021) discover that the optimization speed varies across different modalities. Therefore, when optimizing a multimodal model with a unified objective, the weak modality fails to learn adequately, leading to modality laziness. Most methods maintain multimodal data inputs during training, achieving modality balance by adjusting optimization objectives and speeds.
Some try to alleviate modality laziness by enforcing the weak modality through modality-specific optimization objectives Wang et al. (2020); Du et al. (2021); Xu et al. (2023); Du et al. (2023); Fan et al. (2023). Wang et al. (2020) propose Gradient-Blending, which calculates the optimal mixing mode of modality losses by determining the overfitting situation for each modality. Du et al. (2021) attempt to distill knowledge from well-trained unimodal models to enhance the unimodal encoders. Fan et al. (2023) introduce the prototype cross-entropy loss for each modality to accelerate the slow-learning modality. Others try to align optimization speed by changing learning rates (Sun et al., 2021) or modifying gradients (Peng et al., 2022; Li et al., 2023; Sun et al., 2023; Fu et al., 2023; Kontras et al., 2024) according to unimodal performance. Sun et al. (2021) dynamically adjust the learning rates of different modalities based on the unimodal predictive loss. Peng et al. (2022) adaptively modulate the gradients of each modality by monitoring the discrepancy in their contribution to the learning objective. These methods promote modality balance and enhance the expressive power of multimodal models. However, they fail to recognize that even when modality laziness is addressed, modality clash persists, meaning that multimodal capability remains limited.
@@ 2.2 Balance with Unimodal Single Input
Some methods attempt to use alternating or selective unimodal inputs to promote multimodal learning, thereby avoiding interference from other modalities and enhancing the expressive power of multimodal models (Wu et al., 2022; Zhang et al., 2024b; Wei et al., 2024; Zhou et al., 2023). Wu et al. (2022) measure the relative speed of each modality and train unimodal branches to fully utilize them. Zhang et al. (2024b) decompose multimodal learning into alternating unimodal learning with gradient modification to preserve cross-modal interactions. Wei et al. (2024) design a sample-level modality contribution evaluation based on Shapley values and perform resampling of the weak modality to enhance specific modality training. These methods, which rely on unimodal inputs, seem to avoid modality clash. However, upon closer inspection, we find that modality clash is introduced at the batch level. Like the methods in Section 2.1, they have not analyzed or addressed the root cause of modality clash.
Overall, current work focuses on addressing modality laziness but fails to recognize that modality clash still exists in balanced multimodal learning. Meanwhile, some methods are constrained by specific model architectures (Wu et al., 2022; Zhang et al., 2024b) or hinder training efficiency (Zhang et al., 2024b; Wei et al., 2024), limiting their applicability. In this paper, we aim to simultaneously mitigate modality laziness and modality clash while ensuring compatibility with various fusion methods and model architectures.
@@ 3.1 Model formulation
In this paper, we focus on the multimodal discrimination task, following related works (Peng et al., 2022; Fan et al., 2023; Wei et al., 2024; Zhang et al., 2024b). For convenience, we consider two input modalities: $m_{a}$ and $m_{v}$. The training dataset is $\mathcal{D}=\{x_{i},y_{i}\}_{i=1,2,\ldots,N}$. Each $x_{i}$ consists of multimodal inputs, i.e., $x_{i}=(x_{i}^{a},x_{i}^{v})$. $y_{i}\in\{1,2,\ldots,M\}$, where $M$ is the number of classes. The multimodal model is trained using batch gradient descent, with data $\boldsymbol{x}=(\boldsymbol{x}^{a},\boldsymbol{x}^{v})$ sampled from a batch $B$.
We use a multimodal model consisting of two unimodal branches for prediction. Each branch has a unimodal encoder, denoted as $\phi^{a}$ and $\phi^{v}$, used to extract features from the corresponding modality of $\boldsymbol{x}$. The encoder outputs are represented as $\boldsymbol{z}^{a}=\phi^{a}(\theta^{a},\boldsymbol{x}^{a})$ and $\boldsymbol{z}^{v}=\phi^{v}(\theta^{v},\boldsymbol{x}^{v})$, where $\theta^{a}$ and $\theta^{v}$ are the parameters of the encoders. The results of the two unimodal encoders are fused in some way (Owens & Efros, 2018; Gunes & Piccardi, 2005) to obtain the multimodal output. We use cross-entropy (CE) loss as the loss function. To achieve better representation ability in the warm-up stage, we add a separate classification head for each modality, as Wang et al. (2020) do, and modify the loss function as follows:
@@ 3.2 Data Remixing Method
Decouple Multimodal Data. In multimodal learning, the model tends to fit the modality that optimizes faster, often converging before the other modality has been sufficiently learned (Du et al., 2021). Therefore, appropriately reducing the training speed of the strong modality may be an effective strategy for balancing learning (Peng et al., 2022). However, modality-level control over the optimization process or objectives overlooks the imbalance at the sample level between modalities (Wei et al., 2024), which limits the effectiveness of the balancing. Based on this premise, we consider promoting balance by assigning specific samples (in terms of difficulty and quantity) to different modalities, achieving a more precise and convenient modality balance.
We first define a sample-level unimodal capability evaluation method to achieve either-or data allocation, instead of using Shapley values as proposed by Wei et al. (2024) [^1] For $x_{i}$, if the unimodal predictions are correct with probabilities of 0.9 and 0.6, the Shapley values for both modalities would be the same, but the representational abilities are different., to avoid dataset expansion. For each sample $x_{i}$, we obtain unimodal prediction probabilities $p_{i}^{k}$, $k\in\{a,v\}$. The representational ability of each modality is then assessed by calculating the KL divergence between $p_{i}^{k}$ and a uniform distribution as
@@
where $p^{k}_{i,j}$ represents the predicted probability for class $j$, and $u_{j}=\frac{1}{M}$ corresponds to the uniform distribution. KL divergence measures the difference between two probability distributions. In the current application scenario, a smaller KL divergence indicates that the output is closer to a uniform distribution, meaning that the separability of the corresponding unimodal output is worse. This can be interpreted as insufficient training for the current modality on this sample. Therefore, we decouple the multimodal inputs for each sample and only retain the modality that performs the worst, masking input from other modalities to achieve accurate unimodal training (Zhang et al., 2024a; Wei et al., 2024). That is, the new training set is as follows:
@@
Without loss of generality, we assume that the multimodal fusion method selected for the model is Concatenation, as done by others (Fan et al., 2023; Peng et al., 2022; Wei et al., 2024). Let $\boldsymbol{W}\in\mathbb{R}^{M\times(d_{z^{a}}+d_{z^{v}})}$ and $\boldsymbol{b}\in\mathbb{R}^{M}$ represent the parameters of the linear classifier that produces the logits output. We can then express the output of the multimodal model (without softmax) as
@@ 4.1 Dataset and Experimental settings
CREMA-D (Cao et al., 2014) is an audiovisual dataset for emotion recognition, consisting of 7,442 video clips from 91 actors. The dataset includes six of the most common emotions: anger, disgust, fear, happy, neutral and sad. A total of 2,443 participants rated each clip for emotion and emotional intensity using three modalities: audiovisual, video only, and audio only. The entire dataset is randomly divided into a training and validation set of 6,698 samples and a test set of 744 samples, with a ratio of approximately 9:1.
Kinetic-Sounds (Arandjelovic & Zisserman, 2017) is a dataset derived from the Kinetics dataset (Kay et al., 2017), which includes 400 action classes based on YouTube videos. Kinetic-Sounds consists of 31 action categories, selected for their potential to be represented both visually and aurally, such as playing various instruments. Each video is manually annotated for human actions using Mechanical Turk and is cropped to 10 seconds, focusing on the action itself. The dataset comprises 19k 10-second video clips, split into 15k for training, 1.9k for validation, and 1.9k for testing.
Experimental settings. Unless otherwise specified, all feature extraction networks used in the experiments are ResNet-18 (He et al., 2016), trained from scratch. During training, we use the Adam (Kingma, 2014) optimizer with $\beta=(0.9,0.999)$ and set the learning rate to $5\times 10^{-5}$. We obtain unimodal prediction results from the unimodal classification heads and also present the results of our method by masking specific modalities to zero, as in Hinton (2012) and Wei et al. (2024). All reported results are averages from three random seeds, with all models trained on two NVIDIA RTX 3090 GPUs using a batch size of 64.
@@ 4.2 Comparison with conventional fusion methods
We first compare our method with several representative multimodal fusion methods commonly used in deep learning frameworks: Concatenation (Concat) (Owens & Efros, 2018), Summation (Sum), Decision Fusion (Decision) (Gunes & Piccardi, 2005), FiLM (Perez et al., 2018), and Bi-Gated (Kiela et al., 2018). The results are shown in Table 1. We apply Data Remixing in combination with Concatenation as the representative fusion method for our approach. It is evident that our method significantly improves model performance across different datasets, with an increase of 8.20% on CREMAD and 5.40% on Kinetic-Sounds, outperforming other conventional fusion strategies. To further demonstrate the generalizability of our method, we combine Data Remixing with Summation and Decision Fusion, achieving substantial improvements across both datasets.
@@ 4.3 Comparison with imbalanced multimodal learning methods
Our method is primarily designed to address the issues of modality imbalance and insufficiency in multimodal learning. To evaluate the improvements, we compare our approach with several representative methods that target multimodal balance and sufficiency: G-Blend (Wang et al., 2020), OGM-GE (Peng et al., 2022), Greedy (Wu et al., 2022), PMR (Fan et al., 2023), MLA (Zhang et al., 2024b) and Resample (Wei et al., 2024). For a fair comparison, we adopt Concatenation as the fusion strategy for the baseline, following the approach used in the above methods. The results, including accuracy (Acc) and dataset expansion factor (Factor), are presented for all methods.
@@ 4.4 Combination with complex cross-modal architectures
The above methods and experiments are conducted using simple fusion methods, where multimodal fusion occurs after the unimodal encoders or classifiers. To validate the applicability of the Data Remixing method in more complex multimodal scenarios, we combine it with two intermediate fusion methods: MMTM (Joze et al., 2020) and CentralNet (Vielzeuf et al., 2018). MMTM recalibrates the channel features of different CNN streams through squeezing and multimodal excitation steps, while CentralNet uses unimodal hidden representations alongside a central joint representation at each layer, performing fusion through a weighted summation learned during training.
@@ 4.5.2 Study of Decouple Multimodal Data
In the process of conducting sample-level evaluation and decoupling multimodal data, we choose to mask the better-performing modality and train with specific samples to alleviate modality imbalance. Figure 3 shows the distribution of samples assigned to different modalities during training. It is clear that the strong modality consistently receives fewer samples compared to the weak modality. Figure 3 illustrates the change in imbalance ratio $\rho$ (Peng et al., 2022) before and after decoupling the modality inputs. The results indicate that our method alleviates the modality imbalance, supporting our hypothesis that balancing modality representation via sample quantity and difficulty can improve model performance. Note that our method is applied after a 10-epoch warm-up stage.